OpenStudy (anonymous):

Find gof if f(x)=x+(1/x) and g(x)= (x+1)/(x+2)

OpenStudy (princeharryyy):

g(f(x))= (f(x)+1)/(f(x)+2)

OpenStudy (princeharryyy):

just out the value of f(x) in place of f(x) in above equation n solve

OpenStudy (princeharryyy):

@Theanitrix

OpenStudy (princeharryyy):

put*

OpenStudy (anonymous):

I know that, I just am not sure how to solve it

OpenStudy (princeharryyy):

ok

OpenStudy (anonymous):

the fraction confuses me

OpenStudy (jhannybean):

\[(g \circ f) = g\left(x+\frac{1}{x}\right) = \frac{\left(x+\dfrac{1}{x}\right)+1}{\left(x+\dfrac{1}{x}\right)+2}\]

OpenStudy (princeharryyy):

well, you got the answer now, I guess.

OpenStudy (jhannybean):

That isn't the answer.

OpenStudy (princeharryyy):

that's almost the final answer @Jhannybean

OpenStudy (anonymous):

the answer would be (x^2+x+1)/(x+1)^2

OpenStudy (anonymous):

but i want to know how they got that answer

OpenStudy (jhannybean):

multiply both the numerator and denominator by \(x\). \[\frac{\left(x+\dfrac{1}{x}\right)+1}{\left(x+\dfrac{1}{x}\right)+2} \cdot \frac{x}{x} \]

OpenStudy (princeharryyy):

(g∘f)=g(x+1/x)=(x^2+1+x)/(x^2+2+x)

OpenStudy (anonymous):

is that because it is the denominator of 1/x? @Jhannybean

OpenStudy (jhannybean):

Yes @Theanitrix

OpenStudy (anonymous):

okay so would that same thing apply to fof?

OpenStudy (jhannybean):

\[(f\circ f)= f\left(x+\frac{1}{x}\right) = \left(x+\frac{1}{x}\right)+\dfrac{1}{x+\dfrac{1}{x}}\]

OpenStudy (jhannybean):

but when you're trying to solve this... set all that complicated junk = \(a\) or any other variable. find the common denominator, and simplify.

OpenStudy (jhannybean):

\[\text{let a}=x+\frac{1}{x}\]\[(f\circ f) = a+\frac{1}{a} = \frac{a^2+1}{a} = \dfrac{\left(x+\dfrac{1}{x}\right)^2+1}{\left(x+\dfrac{1}{x}\right)} \]

OpenStudy (jhannybean):

\[=\frac{x^2+\dfrac{1}{x^2}+2 +1}{\dfrac{x^2+1}{x}} =\frac{x^2+\dfrac{1}{x^2}+3}{\dfrac{x^2+1}{x}} \] So now you'd want to multiply both numerator and denominator by \(\dfrac{x}{x}\)

OpenStudy (anonymous):

i did it a different way

OpenStudy (anonymous):

and got the answer x^4+3x^2+1/x(x^2+1)

OpenStudy (jhannybean):

Thats correct

OpenStudy (anonymous):

okay, i understood your approach as well so thank you

OpenStudy (jhannybean):

\[\frac{x^2+\dfrac{1}{x^2}+3}{\dfrac{x^2+1}{x}} \cdot \frac{x}{x} =\frac{\dfrac{x^4+3x^2+1}{x}}{x^2+1} = \color{red}{\frac{x^4+3x^2+1}{x(x^2+1)}}\]